OpenStudy (anonymous):

Solve. 2x2 = 8 x = 2 x = 2 and x = –2 x = –2 and x = 4 x = 4

OpenStudy (anonymous):

do you mean \[2x ^{2}=8\]

OpenStudy (anonymous):

x=+,-2

OpenStudy (anonymous):

yes im sorry @MathDad72

OpenStudy (anonymous):

that's ok you want to go through the steps to solve? or just use the answers given?

OpenStudy (anonymous):

if you could just explain how you got your answer that would be great

OpenStudy (anonymous):

We want to get the x by itself, correct?

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

how would you "move" the 2 over to the other side?

OpenStudy (anonymous):

\[(\frac{ 1 }{ 2 }) 2x ^{2} = 8(\frac{ 1 }{ 2 })\] Does this make sense?

OpenStudy (anonymous):

divide?

OpenStudy (anonymous):

correct if you divide the left side by 2 you will get \[x ^{2}\]

OpenStudy (anonymous):

and what ever you do to one side you need to do to the other (right) \[x ^{2} = 4\]

OpenStudy (anonymous):

still good?

OpenStudy (anonymous):

so now to "move" the square to the other side.

OpenStudy (anonymous):

\[\sqrt{x ^{2}} = \pm \sqrt{4}\]

OpenStudy (anonymous):

\[x = \pm 2\]

OpenStudy (anonymous):

Anything you want to go over in more detail? Or are you good?